OpenStudy (anonymous):

Mark simplified three over five divided by five over eight.; his work is shown below. Identify where he made his error. Original division problem: three over five divided by five over eight. Step 1: three over five multiplied by eight over five. Step 2: three times eight over five times five. Step 3: twenty-four over five. Step 4: four and four over five.

OpenStudy (anonymous):

@mathway @King.Void. @welshfella are any of you available to help me?

OpenStudy (anonymous):

\[ \huge \frac{ \frac{ 3 }{ 5 } }{ \frac{ 5 }{ 8 } }\]

OpenStudy (anonymous):

Based on his work, Step 1 is \[\huge \frac{ 3 }{ 5 } \times \frac{ 8 }{ 5 }\] Is this correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Right, because when you divide, you multiply 3/5 to the reciprocal of 5/8, which is 8/5. :)

OpenStudy (anonymous):

Step 2 is \[\huge \frac{ 3 \times 8}{ 5 \times 5 }\] Is this right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

@mathway

OpenStudy (anonymous):

For Step 3, he multiplied\[ \huge \frac{ 3 \times 8}{ 5 \times 5 }=\frac{ 24 }{ 5 }\]. Is this correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What is 3 x 8?

OpenStudy (anonymous):

What is 5 x 5?

OpenStudy (anonymous):

oh no 3x8= 24 5x5=25

OpenStudy (anonymous):

Yes, so was his Step 3 right or wrong? :)

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

So you can conclude that he made an error in Step 3. :)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thank you (;